OpenStudy (anonymous):

the scale factor of 125% as a fraction and a decimal

OpenStudy (jack1):

In the equation y=Cx, C is the scale factor for x. 125% = 1.25 as percent literally means "per 100" so divide by 100 to get decimal 125% = 5/4 as 1/4 = 25% so 5/4 = ...125%